during an investigation a scientist heated 2.76 g of silver carbonate till it decomposed to leave a silver residue. the total mass of silver residue formed was 2.16 grams. does the law of conservation of mass hold true in this case?

Grade:8

1 Answers

Khimraj
5 years ago
The equation that shows the decomposition of silver carbonate is as follows:
2Ag2CO3 .........> 4Ag + 2CO2 + O2
From the periodic table:
mass of silver = 107.8682 grams
mass of carbon = 12 grams
mass of oxygen = 16 grams
molar mass of Ag2CO3 = 2(107.8682) + 12 + 3(16) = 275.7364 grams
From the balanced equation above:
2(275.7362) = 551.4728 grams of Ag2CO3 produces 4(107.8682) = 431.4728 grams of Ag
Therefore, to know the mass of Ag produced from 2.76 grams of Ag2CO3, we will simply use cross multiplication as follows:
mass of Ag produced = (2.76*431.4728) / (551.4728) = 2.16 grams
Based on the above calculations, the law of conservation of mass is applied.